solve ln(x3) + ln(x^4) = 14
Answer by Alan3354(69443)

ln(x3) + ln(x^4) = 14
---
Is 3 is the exponent of x?
ln(x^3) + ln(x^4) = 14
ln(x^7) = 14
7*ln(x) = 14
ln(x) = 2
x = e^2